A vehicle of mass m is moving on a rough horizontal road with momentum P. If the coefficient of friction between the tyre and road is µ, then the stopping distance is?

Grade:9

1 Answers

Aditya
7 years ago
Initial Velocity=p/m
Final velocity=0(it has to stop)
Acceleration=  --umg/m=-ug
By using  v^2-u^2=2as
By putting value we get
p^2/m^2=2ugs
↔s=p^2/2m^2ug